Shift report 20011114 08:00-16:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Wed Nov 14 2001 - 07:57:25 EST

  • Next message: Michael Murray: "Re: Fw: RHIC Au x Au Running at 22 GeV"

    ------------------------------------------------------------
    Date: Wed Nov 14 07:57:25 2001, Shift: 08:00-16:00
    Supervisor: TML, with: 
    Log book#10 pages 137 - 138
    
    Shift Summary:
    ------------------------------------------------------------
    1. The machine is up and running. 2 good stores since 0300.
    First ended by QLI. Next one came 1hour and 15 min later.
    
    2 run    trig1   trig6   
      5790   274k    87k   
      5791   321k    103k  
      5792   still running  moved MRS @ 40deg 4B, FFS @ 12deg .5A
    
    3. Have not moved the BFS to 12 degrees since the inner 
    camera is notworking. DB is going to request an access.
    Otherwise no problems.
    
    
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Wed Nov 14 2001 - 07:58:00 EST